Find the equation of the line with x-intercept 1 and y-intercept 5.
Answer by mananth(16946)

The coordinates of intersection are (1,0) & (0,5)
x1 y1 x2 y2
1 0 0 5

slope m = (y2-y1)/(x2-x1)
( 5 - 0 )/( 0 - 1 )
( 5 / -1 )
m= -5
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( 1 , 0 ) in
Y = m x + b
0.00 = -5 * 1 + b
0.00 = -5 + b
b= 0 + 5
b= -5
So the equation will be
Y = -5 x + 5
m.ananth@hotmail.ca